# A causal viewpoint on prediction model performance under changes in case-mix: discrimination and calibration respond differently for prognosis and diagnosis predictions

### v2

@@ Abstract
Prediction models inform important clinical decisions, aiding in diagnosis, prognosis, and treatment planning. The predictive performance of these models is typically assessed through discrimination and calibration. However, changes in the distribution of the data impact model performance. In health-care, a typical change is a shift in case-mix: for example, for cardiovascular risk management, a general practitioner sees a different mix of patients than a specialist in a tertiary hospital.
This work introduces a novel framework that differentiates the effects of case-mix shifts on discrimination and calibration based on the causal direction of the prediction task. When prediction is in the causal direction (often the case for prognosis predictions), calibration remains stable under case-mix shifts, while discrimination does not. Conversely, when predicting in the anti-causal direction (often with diagnosis predictions), discrimination remains stable, but calibration does not.
A simulation study and empirical validation using cardiovascular disease prediction models demonstrate the implications of this framework. This framework provides critical insights for evaluating and deploying prediction models across different clinical settings, emphasizing the importance of understanding the causal structure of the prediction task.
@@ 1 Introduction
An issue with predictive performance is that there may be important changes between when a model’s predictive performance was last evaluated, and when and where it is used, meaning that the underlying data distribution may have changed. No model can have good predictive performance under all arbitrary changes in the data distribution, but we may consider one important class of changes in distribution described with the term ‘case-mix’. For instance, when comparing the general practitioner (GP) setting with a tertiary hospital setting, the frequency of certain comorbidities and risk factors will be different across settings. Typically the tertiary center will encounter more high risk patients, so their ‘case-mix’ is different than in the GP setting. Another change in case-mix is the frequency of myocardial infarction in patients presenting with chest pain at either the GP or in those referred to acute cardiac care centers.
@@
Changes in the data distribution necessarily lead to changes in predictive performance. However, we introduce a new framework that shows that when the prediction task is in the causal direction, which is often the case in prognosis predictions, versus in the anti-causal direction, often the case when predicting a diagnosis, a change in case-mix has a different interpretation. Importantly, depending on the prediction direction, either calibration or discrimination is preserved under shifts in case-mix, but never both. We achieve this by looking at diagnosis and prognosis through a causal lens. Inferring a diagnosis is typically prediction in the anti-causal direction, meaning predicting the cause (=underlying diagnosis) based on its effects (=symptoms), and here a change in case-mix is a change in distribution of the prediction target (the diagnosis). In contrast, prognosis is typically prediction in the causal direction, meaning a future outcome predicted from current patient characteristics, and here a shift in case-mix is a change in the distribution of patient characteristics. The crucial insight of our framework is that a prediction model’s discrimination depends on the distribution of the features given the outcome ($X$ given $Y$) and is thereby invariant to changes in the marginal distribution of the outcome. Conversely, calibration depends on the distribution of the outcome given the features ($Y$ given $X$) and is thereby invariant to changes in the marginal distribution of the features. See Figure 1 for a schematic overview.
@@
Our new perspective sheds light on changes in predictive performance under shifts in case-mix. Our result shows that the causal direction of the prediction has important implications for the evaluation and deployment of prediction models. For example, when evaluating a model used for prognosis across different settings, changes in discrimination are expected under shifts in case-mix, but changes in calibration are not, and vice-versa for diagnostic models. When re-evaluating a prognostic model in a different setting, a change in discrimination is expected and thus no cause for concern. However, a marked change in calibration may warrant further research. This perspective helps developers and guideline makers judge where and when a prediction model has dependable predictive performance. Also, depending on the task and whether discrimination or calibration is more important, prediction model developers may improve the robustness of their model to changes in case-mix by only including variables in the prediction model that are either all causal or all anti-causal but not mixed when possible.
@@
To introduce the framework, we first review the concepts of discrimination and calibration and then we define changes in case-mix from a causal viewpoint. Next put the two pieces together in a new framework and answer: when to expect what changes in predictive performance? We illustrate the result with a simulation study and test the framework empirically in a systematic review of 1382 prediction models, where we find that prognostic models indeed have more variance in discrimination when tested in external validation studies. Finally we discuss how this theory can be applied in practice.
2 Notation and Recap of predictive performance: discrimination and calibration
We consider prediction models of a binary outcome $Y$ using features $X$ with a prediction model $f:\mathcal{X}\to[0,1]$. The features can come from an arbitrary (multi-)dimensional distribution but we assume the model’s predictions $f(X)$ can reasonably be considered as continuous. We will denote environments with an environment variable $E$ where for example $E=0$ may be a GP setting and $E=1$ a community hospital, $E=2$ a university medical center and so on[bareinboimCausalInferenceDatafusion2016]. With $P(.)$ we will denote (conditional) distributions or densities over random variables, for example $P(Y|X)$ denotes the distribution of outcome $Y$ given features $X$.
@@ 2.1 Discrimination: sensitivity, specificity and AUC
Typical metrics of discrimination are sensitivity (sometimes called recall), specificity and AUC. Sensitivity is the ratio of true positives over the total number of positive cases. Specificity is the ratio of true negatives over the total number of negative cases. To calculate sensitivity and specificity, we need to choose a threshold $10<\tau<$ 1 for the output of $f(X)$ and label all $f(X)>\tau$ as positive predicted cases and $f(X)\leq\tau$ as negative predicted cases. This results in a 2 by 2 table with predicted versus actual outcomes (sometimes called the ‘confusion table’), see Table 1. By varying $\tau$ between 0 and 1 we get a range of values for specificity and specificity. Plotting these in the receiver-operating-curve and calculating the area under this curve we get the popular AUC metric or c-statistic. Note that for calculating sensitivity we only need the positive cases ($Y=1$), and for specificity we only need the negative cases ($Y=0$). Measures of discrimination depend on the distribution of the prediction (and thus the features) given the outcome. This immediately implies that if we were to only change the ratio of positive and negative cases through some hypothetical intervention, the sensitivity and specificity will remain unchanged, and thus the resulting AUC. Therefore it is sometimes said that sensitivity and specificity are prevalence independent.
@@
Table 1: Confusion table. By specifying a threshold $10<\tau<$ 1 for a prediction model $f:\mathcal{X}\to[0,1]$ and tabulating the results against the ground truth outcome $Y\in\{0,1\}$, one gets the confusion table and can calculate metrics of discrimination such a sensitivity and specificity.
@@ 3 A causal framework for predictive performance under changes in case-mix
Since discrimination depends on the distribution of the features given the outcome ($X$ given $Y$) but calibration on the distribution of the outcome given the features ($Y$ given $X$), we may expect metrics of discrimination and calibration to respond differently when changes occur in the marginal distribution of $X$ or $Y$. In this section we first formalize the notion of a shift in case-mix and how this depends on whether a prediction is in the causal direction (future outcome given features) or the anti-causal direction (disease given symptoms). Then we will draw the connection between the two insights leading to our main result.
@@ 3.1 A shift in case-mix is a change in the marginal distribution of the cause variable
We define a shift in case-mix between different environments (e.g. GP versus hospital setting) as a change in the marginal distribution of the cause variable. When the prediction is in the causal direction, a shift in case-mix is a change in the marginal distribution of the features $X$, whereas when the prediction is in the anti-causal direction it is a shift in the marginal distribution of the outcome $Y$. Finally, the prediction problem could be neither causal or anti-causal, but confounded by another variable $Z$, in that case the shift is in the distribution of the confounder $Z$. See Figure 2 for directed acyclic graphs (DAGs) depicting these situations and Table 2 for an overview with examples. We give a formal definition in the Appendix 2.
@@
Figure 2: directed acyclic graphs for 2-variable prediction problems with a shift in case-mix, meaning the environment variable only affects the marginal distribution of one variable but not a conditional distribution. The prediction is always made from $X$ to $Y$. $X$ denotes an arbitrary set of features with arbitrary feature types.
Each of the DAGs in Figure 2 encodes different conditional idependencies. Specifically the DAG in the causal direction (Figure 2(a)) implies that $Y$ is independent of $E$ given $X$. This entails that the distribution $P(Y|X)$ is transportable across environments, so for different environments $E=0,1,\ldots$, $P(Y|X,E=0)=P(Y|X,E=1)=P(Y|X)$, whereas in the anti-causal direction (Figure 2(b)) the distribution $P(Y|X)$ is transportable, meaning $P(X|Y,E)=P(X|Y)$ [bareinboimCausalInferenceDatafusion2016]. In the confounded DAG (Figure 2(c)) neither $P(Y|X)$ or $P(X|Y)$ is transportable.
In the DAGs in Figure 2 the environment variable influences the cause variable ($X,Y$ or $Z$) but not the effect variable ($Y$ or $X$). Why exclude arrows from the environment to the effect variable in the definition of a shift in case-mix? If there is an arrow from environment to the effect variable, neither $P(Y|X)$ or $P(X|Y)$ are transportable across environments so nothing can be said regarding the calibration and discrimination of a prediction model on an unseen environment based on data from the observed environments only. Also, based on concrete clinical settings it may be reasonable based on patient selection mechanisms to assume that at least the distribution of the cause variable differs between environments, but maybe not the effect given the cause. See the Appendix 5.1 for several concrete medical examples where these assumptions may hold and where they are violated.
@@
When predicting in the anti-causal direction (often with diagnosis predictions), a shift in case-mix across environments means a shift in the marginal distribution of the outcome, and discrimination remains stable but not calibration. Conversely, when predicting in the causal direction (often with prognosis predictions), a shift in case-mix across environments means a shift in marginal distribution of the features, and calibration remains stable, but not discrimination.
When $f$ is perfectly calibrated on an environment, it will remain perfectly calibrated under shifts of the marginal distribution of the features (see Theorem 1 in the Appendix). Note that when $f$ is not perfectly calibrated and this mis-calibration depends on $X$, in general the average calibration will also change when predicting in the causal direction.
An important implication of this result is that discrimination or calibration may be preserved under changes in case-mix, but never both
As a remark, we note that perfectly calibrated models obviously cannot be better calibrated in other environments, so any change in calibration necessarily implies a worsening of calibration. For discrimination, this is not automatically the case. In fact, models can show better discrimination in other environments when the distribution of outcome probabilities is less concentrated around 50%.
@@ 4.1 Illustrative simulation
We evaluate both models in three hypothetical environments: a screening environment (with low outcome prevalence), a GP setting (with intermediate prevalence) and a hospital setting (high prevalence). For the prognosis model, the marginal distribution of $X$ depends on the environment through $\alpha_{e},\beta_{e}$, but not the distribution of $Y$ given $X$. For the diagnosis model, the marginal distribution of $Y$ depends on the environment through $P_{e}$, but not the distribution of $X$ given $Y$. The different values for these parameters are given in Table 3
task parameter screening GP hospital
@@
Table 3: Values for different simulation parameters in three hypothetical environments. GP = general practitioner
In Figure 3 we show the results of training a prediction model in the screening environment and evaluating it either in the same environment (=internal validation) or in a different environment (=external validation). For the prognostic model the calibration remains the same across environments (though some values of $P(Y=1|X)$ become very rare because of the shift in distribution of $X$). The discrimination changes across environments. For the diagnostic model, the reverse is true.
Figure 3: Overview figure of illustrative simulation experiment of a model trained on data from a screening environment, and evaluated on either the screening environment (=internal validation) or the GP environment or the hospital environment (=external validation), with increasing outcome probabilities. For models predicting in the anti-causal direction (e.g. diagnostic models), a shift in case-mix entails an intervention on the distribution of the predicted outcome, so discrimination remains the same but calibration changes. For models predicting in the causal direction (e.g. prognosis models), a shift in case-mix entails an intervention on the distribution of the features, so calibration remains the same but the discrimination changes. The discrimination facets are receiver-operating-curves with on the horizontal axis 1 minus specificity and on the vertical axis sensitivity. The calibration facets are calibration curves with on the horizontal axis the predicted probability and on the vertical axis the actual probability.
By repeating this process for each of the three environments, each time training on one environment and evaluating on all environments for both the causal prediction model and the anti-causal model, we get in total six models, each evaluated three times. We measure discrimination with AUC for discrimination, and calibration error as the absolute difference between the predicted outcome probability and the actual outcome probability for each observation: $\frac{1}{N}\sum_{i}^{N}\left|P(Y=1|X=x_{i})-f(x_{i})\right|$ (analogous to the Integrated Calibration Error defined in [austinIntegratedCalibrationIndex2019]). Plotting these 18 points on 6 lines in 2 dimensions leads to an interesting pattern, where the models predicting in the causal direction are easily discernible from those predicting in the anti-causal direction (Figure 4). In the Appendix 5.2 we provide visualizations of $P(Y|X)$ and $P(X|Y)$ for the different environments and tasks.

### v3

@@ Abstract
Prediction models need reliable predictive performance as they inform clinical decisions, aiding in diagnosis, prognosis, and treatment planning. The predictive performance of these models is typically assessed through discrimination and calibration. Changes in the distribution of the data impact model performance and there may be important changes between a model’s current application and when and where its performance was last evaluated. In health-care, a typical change is a shift in case-mix. For example, for cardiovascular risk management, a general practitioner sees a different mix of patients than a specialist in a tertiary hospital.
This work introduces a novel framework that differentiates the effects of case-mix shifts on discrimination and calibration based on the causal direction of the prediction task. When prediction is in the causal direction (often the case for prognosis preditions), calibration remains stable under case-mix shifts, while discrimination does not. Conversely, when predicting in the anti-causal direction (often with diagnosis predictions), discrimination remains stable, but calibration does not.
A simulation study and empirical validation using cardiovascular disease prediction models demonstrate the implications of this framework. The causal case-mix framework provides insights for developing, evaluating and deploying prediction models across different clinical settings, emphasizing the importance of understanding the causal structure of the prediction task.
keywords: calibration, discrimination, case-mix, causal inference, prediction model, external validation
@@ 1 Introduction
An issue with predictive performance is that there may be important changes between when a model’s predictive performance was last evaluated, and when and where it is used, meaning that the underlying data distribution may have changed. No model can have good predictive performance under all arbitrary changes in the data distribution, but we may consider one important class of changes in distribution described with the term ‘case-mix’. For instance, when comparing cardiovascular risk management in the general practitioner setting with a tertiary hospital setting, the frequency of certain comorbidities and risk factors will be different across settings. Typically the tertiary center will encounter more high risk patients, so their ‘case-mix’ is different than in the general practitioner setting. Another change in case-mix is the frequency of myocardial infarction in patients presenting with chest pain at either the general practitioner or in those referred to acute cardiac care centers. We present a formal definition of a shift in case-mix using the language of causality. The independent causal mechanisms principle states that when viewing the joint distribution of observed data trough a mechanistic causal lens, where effects are created from inputs (causes) by a causal mechanism, the distribution of the inputs (causes) is independent of the mechanism that produces the outputs from the inputs [petersElementsCausalInference2017]. A natural causal definition of a shift in case-mix is thus a shift in the marginal distribution of the causes, and we may make the additional assumption that the mechanism is the same across environments. When applied to the above example this would mean that though tertiary case center patients have a different distribution of risk factors than patients from the general practitioner setting, one may hope that given knowledge of sufficient risk-factors, the risk of cardiovascular disease for two patients with the same values of risk-factors is the same regardless of what setting they are in.
@@
This causal definition of a shift in case-mix implies that when the prediction task is in the causal direction versus in the anti-causal direction, a change in case-mix has a different interpretation. Inferring a diagnosis is typically prediction in the anti-causal direction, meaning predicting the cause (=underlying diagnosis) based on its effects (=symptoms), and here a change in case-mix is a change in distribution of the prediction target (the diagnosis). In contrast, prognosis is typically prediction in the causal direction, meaning a future outcome predicted from current patient characteristics, and here a shift in case-mix is a change in the distribution of patient characteristics. Importantly, depending on the prediction direction, either calibration or discrimination is preserved under shifts in case-mix, but not both. The crucial insight underlying our results is that a prediction model’s discrimination depends on the distribution of the features given the outcome ($X$ given $Y$) and is thereby invariant to changes in the marginal distribution of the outcome. Conversely, calibration depends on the distribution of the outcome given the features ($Y$ given $X$) and is thereby invariant to changes in the marginal distribution of the features. See Figure 1 for a schematic overview.
@@
Our result shows that the causal direction of the prediction has important implications for the development, evaluation and deployment of prediction models. For example, when evaluating a model used for prognosis across different settings, changes in discrimination are expected under shifts in case-mix, but changes in calibration are not, and vice-versa for diagnostic models. When re-evaluating a prognostic model in a different setting, a change in discrimination is expected and thus no cause for concern. However, a marked change in calibration may warrant further research. Another perhaps unexpected result is that when a model is evaluated across different environments, the observation that either discriminaton or calibration remains stable is a stronger sign of robustness to changes in environment than when both remain stable. The reason is that when both remain stable, this proves that the testing environments were not meaningfully different. Only when either discrimination or calibration changes and the other is stable, we gain some confidence that the model remains robust across different environments. This perspective helps developers and guideline makers judge where and when a prediction model has dependable predictive performance. Also, depending on the task and whether discrimination or calibration is more important, prediction model developers may improve the robustness of their model to changes in case-mix by only including variables in the prediction model that are either all causal or all anti-causal but not mixed, when possible.
@@
To introduce the framework, we first review the concepts of discrimination and calibration and then we define changes in case-mix from a causal viewpoint. Next put the two pieces together in a new framework and answer: when to expect what changes in predictive performance? We illustrate the result with a simulation study and test the framework empirically in a systematic review of 1382 prediction models, where we find that prognostic models indeed have more variance in discrimination when tested in external validation studies. Finally we discuss how this theory can be used in practice.
2 Notation and review of predictive performance: discrimination and calibration
We consider prediction models of a binary outcome $Y$ using features $X$ with a prediction model $f:\mathcal{X}\to[0,1]$. The features can come from an arbitrary (multi-)dimensional distribution.We will denote environments with an environment variable $E$ where for example $E=0$ may be a general practitioner setting, $E=1$ a community hospital and $E=2$ a university medical center [bareinboimCausalInferenceDatafusion2016]. With $P(.)$ we will denote (conditional) distributions or densities over random variables, for example $P(Y|X)$ denotes the distribution of outcome $Y$ given features $X$.
@@ 2.1 Discrimination: sensitivity, specificity and AUC
The typical metrics of discrimination are sensitivity (sometimes called recall), specificity and AUC. Sensitivity is the ratio of true positives over the total number of positive cases. Specificity is the ratio of true negatives over the total number of negative cases. To calculate sensitivity and specificity, we need to choose a threshold $0\leq\tau\leq 1$ for the output of $f(X)$ and label all $f(X)>\tau$ as positive predicted cases and $f(X)\leq\tau$ as negative predicted cases. This results in a 2 by 2 table with predicted versus actual outcomes (sometimes called the ‘confusion table’), see Table 1. By varying $\tau$ between 0 and 1 we get a range of values for sensitivity and specificity. Plotting these in the receiver-operating-curve and calculating the area under this curve we get the popular AUC metric or c-statistic. Note that for calculating sensitivity we only need the positive cases ($Y=1$), and for specificity we only need the negative cases ($Y=0$). Measures of discrimination depend on the distribution of the prediction (and thus the features) given the outcome. This immediately implies that if we were to only change the ratio of positive and negative cases through some hypothetical intervention, the sensitivity and specificity will remain unchanged, and thus the resulting AUC. Therefore it is sometimes said that sensitivity and specificity are prevalence independent.
@@
Table 1: Confusion table. By specifying a threshold $0\leq\tau\leq 1$ for a prediction model $f:\mathcal{X}\to[0,1]$ and tabulating the results against the ground truth outcome $Y\in\{0,1\}$, we get the confusion table and can calculate metrics of discrimination such a sensitivity and specificity.
@@ 3 A causal framework for predictive performance under changes in case-mix
Since discrimination depends on the distribution of the features given the outcome ($X$ given $Y$) but calibration on the distribution of the outcome given the features ($Y$ given $X$), we may expect metrics of discrimination and calibration to respond differently when changes occur in the marginal distribution of $X$ or $Y$. In this section we first formalize the notion of a shift in case-mix and how this depends on whether a prediction is in the causal direction (future outcome given features) or the anti-causal direction (e.g. disease given symptoms). Then we will draw the connection between the two insights leading to our main result.
@@ 3.1 A shift in case-mix is a change in the marginal distribution of the cause variable
Inspired by the principle of independence of cause and mechanism [petersElementsCausalInference2017, schoelkopfCausalAnticausalLearning2012], we define a shift in case-mix between different environments (e.g. general practitioner versus hospital setting) as a change in the marginal distribution of the cause variable. When the prediction is in the causal direction, a shift in case-mix is a change in the marginal distribution of the features $X$, whereas when the prediction is in the anti-causal direction it is a shift in the marginal distribution of the outcome $Y$.
Finally, the prediction problem could be neither causal or anti-causal, but confounded by another variable $Z$, in that case the shift is in the distribution of the confounder $Z$. See Figure 2 for directed acyclic graphs (DAGs) depicting these situations and Table 2 for an overview with examples. We give a formal definition in the Appendix 2.
@@
Figure 2: directed acyclic graphs for 2-variable prediction problems with a shift in case-mix, meaning the environment variable only affects the marginal distribution of only the cause variable. The prediction is always made from feature $X$ to outcome $Y$, $E$ denotes the environment.
Each of the DAGs in Figure 2 encodes different conditional idependencies. Specifically the DAG in the causal direction (Figure 2(a)) implies that $Y$ is independent of $E$ given $X$. This entails that the distribution $P(Y|X)$ is transportable across environments, so for different environments $E=0,1,\ldots$, $P(Y|X,E=0)=P(Y|X,E=1)=P(Y|X)$, but $P(X|Y)$ is not transportable: $P(X|Y,E)\neq P(X|Y)$. Conversely, in the anti-causal direction (Figure 2(b)) the distribution $P(X|Y)$ is transportable, meaning $P(X|Y,E)=P(X|Y)$, but not $P(Y|X)$. In the confounded DAG (Figure 2(c)) neither $P(Y|X)$ or $P(X|Y)$ is transportable.
In the DAGs in Figure 2 the environment variable influences the cause variable ($X,Y$ or $Z$) but not the effect variable ($Y$ or $X$). Why exclude arrows from the environment to the effect variable in the definition of a shift in case-mix? First, when viewed as a mechanistic description of the data generating process, the principle of independence of cause and mechanism states that the distribution of the cause variable is independent of the mechanism that produces the effect variable [petersElementsCausalInference2017]. Also, if there is an arrow from environment to the effect variable, neither $P(Y|X)$ or $P(X|Y)$ are transportable across environments so nothing can be said regarding the calibration and discrimination of a prediction model on an unseen environment based on data from the observed environments only and the assumptions expressed in the DAG. Finally, in clinical settings it may be reasonable based on temporal ordering and patient selection mechanisms to assume that at least the distribution of the cause variable differs between environments, but maybe not the effect given the cause. See the Appendix 6.1 for several concrete medical examples where these assumptions may hold.
@@
When predicting in the anti-causal direction (often with diagnosis predictions), a shift in case-mix across environments means a shift in the marginal distribution of the outcome, and discrimination remains stable across environments but not calibration. Conversely, when predicting in the causal direction (often with prognosis predictions), a shift in case-mix across environments means a shift in marginal distribution of the features, and calibration remains stable, but not discrimination.
For prediction in the causal direction, when $f$ is perfectly calibrated on an environment, it will remain perfectly calibrated under shifts of the marginal distribution of the features (see Theorem 1 in the Appendix). Note that when $f$ is not perfectly calibrated and this mis-calibration depends on $X$, in general the average calibration will also change when predicting in the causal direction. An important implication of this result is that discrimination or calibration may be preserved under changes in case-mix, but typically not both [^1] For predicting in the causal direction, examples where the AUC remains constant across environments may be constructed. Consider having a mixture of beta-distributions for $P(Y|X)$ with their modes $\mu_{1}=0.25$ and $\mu_{2}=0.75$. Shifting $\mu_{1}$ to 0 will increase AUC, shifting $\mu_{2}$ to 0.5 will decrease AUC. By shifting both modes at the same time, these changes can be set to balance out..
As a remark, we note that perfectly calibrated models obviously cannot be better calibrated in other environments, so any change in calibration necessarily implies a worsening of calibration. For discrimination, this is not automatically the case. In fact, models show better discrimination in other environments then the training data when the distribution of outcome probabilities becomes less concentrated around 50%.
@@ 4.1 Illustrative simulation
We evaluate both models in three hypothetical environments: a screening environment with low outcome prevalence, a general practitioner setting with intermediate prevalence and a hospital setting with high prevalence. For the prognosis model, the marginal distribution of $X$ depends on the environment through $\alpha_{e},\beta_{e}$, but not the distribution of $Y$ given $X$. For the diagnosis model, the marginal distribution of $Y$ depends on the environment through $P_{e}$, but not the distribution of $X$ given $Y$. The different values for these parameters in the simulation are given in Table 3.
task parameter screening general practitioner hospital
@@
Table 3: Values for different simulation parameters in three hypothetical environments.
In Figure 3 we show the results of training a prediction model in the screening environment and evaluating it either in the same environment (‘internal validation’) or in a different environment (‘external validation’). For the prognostic model the calibration remains the same across environments; the discrimination changes across environments. For the diagnostic model, the reverse is true.
Figure 3: Overview figure of illustrative simulation experiment of a model trained on data from a screening environment, and evaluated on either the screening environment (‘internal validation’) or the general practitioner (GP) environment or the hospital environment (‘external validation’), with increasing outcome probabilities. For models predicting in the anti-causal direction (e.g. diagnostic models), a shift in case-mix entails a shift in the distribution of the outcome, so discrimination remains the same but calibration changes. For models predicting in the causal direction (e.g. prognosis models), a shift in case-mix entails a shift in the distribution of the features, so calibration remains the same but the discrimination changes. The discrimination facets are receiver-operating-curves with on the horizontal axis 1 minus specificity and on the vertical axis sensitivity. The calibration facets are calibration curves with on the horizontal axis the predicted probability and on the vertical axis the actual probability.
By repeating this process for each of the three environments, each time training on one environment and evaluating on all environments for both the causal prediction model and the anti-causal model, we get in total six models, each evaluated three times. We measure discrimination with AUC and calibration error as the average absolute difference between the predicted outcome probability and the actual outcome probability for each observation: $\frac{1}{N}\sum_{i}^{N}\left|P(Y=1|X=x_{i})-f(x_{i})\right|$ (analogous to the Integrated Calibration Error defined in [austinIntegratedCalibrationIndex2019]). Plotting these 18 points on 6 lines in 2 dimensions leads to an interesting pattern, where the models predicting in the causal direction are easily discernible from those predicting in the anti-causal direction (Figure 4). In the Appendix 6.2 we provide visualizations of $P(Y|X)$ and $P(X|Y)$ for the different environments and tasks.